OpenStudy (anonymous):

Simplify (x + 3)(x2 + 4x – 2)

OpenStudy (anonymous):

aka multiply?

OpenStudy (ajprincess):

(x + 3)(x^2 + 4x – 2) =x(x^2+4x-2)+3(x^2 + 4x – 2) =x*x^2+x*4x-x*2+3*x^2+3*4x-3*2 Can u simplify nw?

OpenStudy (anonymous):

first multiply everything in the second parentheses by x you get \(x^3+4x^2-2x\) then multiply all by 3 and get \(3x^2+12x-6\) then combine like terms in \[x^3+4x^2-2x+3x^2+12x-6\] and then tell your teacher that in no sense is the second expression any "simpler" than the first, and the directions should say "multiply" if that is what is meant

OpenStudy (anonymous):

The Royal Fruit Company produces two types of fruit drinks. The first type is 70% pure fruit juice, and the second type is 95% pure fruit juice. The company is attempting to produce a fruit drink that contains 80% pure fruit juice. How many pints of each of the two existing types of drink must be used to make 120 pints of a mixture that is 80% pure fruit juice?

OpenStudy (anonymous):

satellite would you care help?

OpenStudy (anonymous):

Thank you